A model airplane is flying in a circular loop in the vertical plane, with a radius of 80.0 m. What is the minimum speed that the airplane can maintain without being forced out of the loop?


1
Expert's answer
2021-02-09T15:26:12-0500
"F_g=F_c,""mg=\dfrac{mv^2}{R},""v=\sqrt{gR}=\sqrt{9.8\ \dfrac{m}{s^2}\cdot80.0\ m}=28\ \dfrac{m}{s}."